OpenStudy (anonymous):

If a b c and d are in G.P, prove that a-b, b-c , c-d are in geometric prgrogression......

OpenStudy (shubhamsrg):

whats your progress so far ?

OpenStudy (anonymous):

i have to solve it back wards......

OpenStudy (shubhamsrg):

yep.. you'll have to prove (a-b)(c-d)=(b-c)^2 using ac=b^2

OpenStudy (shubhamsrg):

and ad=c^2 ,also ad=bc

OpenStudy (shubhamsrg):

correction: bd=c^2

OpenStudy (shubhamsrg):

its upside down,,please re-adjust..

OpenStudy (shubhamsrg):

anyways,,i could read it,,and you'll done right till now..i'll write the last step here b^2 + c^2 -2bc = ac+bd-bc-ad

OpenStudy (shubhamsrg):

you have to prove LHS = RHS start with any, say LHS

OpenStudy (shubhamsrg):

substitute b^2 = ac c^2 = bd

OpenStudy (shubhamsrg):

also we know bc=ad so 2bc = bc + ad make these substitutions and there it is..

OpenStudy (anonymous):